OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -83?

OpenStudy (anonymous):

i thought a taught you :P

OpenStudy (anonymous):

ar^0+ar^1+ar^2...

OpenStudy (anonymous):

i am getting -3115

OpenStudy (anonymous):

a1=a=13 since a17=-83 then the formula is ar^16 it is 16 because they start at one but our formulas exponents start at 0 -83=13*r^16 sfr do you know how to solve for r?

OpenStudy (anonymous):

Yes and i understood what you meant by it but i do not have the formula to plug the #'s in :"(

OpenStudy (anonymous):

r = e^(ln(b/a)/n) if you wanna solve for r :) in a r^n=b :)

OpenStudy (anonymous):

a=13 and a17=-83. => a+16d=-83 => d=-6 now put this in the formula for the sum S=35/2[2*13+34*-6]=-3115

OpenStudy (anonymous):

-200 -197 -194 -191

OpenStudy (anonymous):

\[An=a1+(n-1)d \]

OpenStudy (anonymous):

and then

OpenStudy (anonymous):

after you find d

OpenStudy (anonymous):

you plug it all into the other equation right?

OpenStudy (anonymous):

a1 = 13 are you sure that is right?

OpenStudy (anonymous):

ooh sorry i am wrong i read that is geometric sequence not arithmetic lol *blush* hold on

OpenStudy (anonymous):

nooooo lol

OpenStudy (anonymous):

ignore all i said and lets start over :P I am so sorry

OpenStudy (anonymous):

tis all good but im strugglin with this topic lol

OpenStudy (anonymous):

-83 = 13 + +(17, -1) d solve for d :)

OpenStudy (anonymous):

d=-6 so use first formula here it is the same formula i use to find d http://en.wikipedia.org/wiki/Arithmetic_progression

OpenStudy (anonymous):

a_n=a_1+(n-1)*d n is ith element an is ith term d is difference and a1 is first term

OpenStudy (anonymous):

13+(35-1)*-6=-191